### v2

# Geolocation with Real Human Gameplay Data: A Large-Scale Dataset and Human-Like Reasoning Framework

Abstract.
Geolocation, the task of identifying an image’s location, requires complex reasoning and is crucial for navigation, monitoring, and cultural preservation. However, current methods often produce coarse, imprecise, and non-interpretable localization. A major challenge lies in the quality and scale of existing geolocation datasets. These datasets are typically small-scale and automatically constructed, leading to noisy data and inconsistent task difficulty, with images that either reveal answers too easily or lack sufficient clues for reliable inference. To address these challenges, we introduce a comprehensive geolocation framework with three key components: GeoComp, a large-scale dataset; GeoCoT, a novel reasoning method; and GeoEval, an aspect-based metric designed to evaluate the correctness of the geolocation reasoning process. At the core of this framework is GeoComp (Geolocation Competition Dataset), a large-scale dataset collected from a geolocation game platform involving 740K users over two years. It comprises 25 million entries of metadata and 3.9 million geo-tagged locations spanning much of the globe, with each location annotated thousands to tens of thousands of times by human users. The dataset offers diverse difficulty levels for detailed analysis and highlights key gaps in current models. Building on this dataset, we propose Geographical Chain-of-Thought (GeoCoT), a multi-step reasoning framework designed to enhance the reasoning capabilities of Large Vision Models (LVMs) in geolocation tasks. GeoCoT improves performance by integrating contextual and spatial cues through a multi-step process that mimics human geolocation reasoning. Finally, we demonstrate that GeoCoT significantly boosts performance by up to 25% on classic geolocation metrics and by 9% in reasoning quality as measured by GeoEval:
Geolocation, Reasoning, Chain of thought, Dataset
copyright: noneisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Natural language generationccs: Computing methodologies Computer vision tasks
1. Introduction
Dataset Size Geographic Source Open Human
Coverage Access Annotation
Google-WS-15k  (Clark et al., 2023a) 15k Global Map Service
GMCP (Zamir and Shah, 2014) 105K Local Map Service
StreetCLIP (Haas et al., 2023) 1M Unknown Map Service
Im2GPS  (Hays and Efros, 2008) 237 Local Web-Scraped
Im2GPS3K  (Vo et al., 2017b) 2997 Local Web-Scraped
YFCC4K  (Vo et al., 2017b) 4536 Local Web-Scraped
YFCC26K  (Theiner et al., 2022a) 26k Local Web-Scraped
MP-16  (Larson et al., 2017) 4.7M Local Web-Scraped
OSV-5M (Astruc et al., 2024) 5.1M Global Map Service
GeoComp 3.3M Global Map Service
Table 1. Comparison of Existing Geolocation Datasets and GeoComp. “Local” refers to city- or region-specific data, while “Global” spans multiple continents. Darker green shades indicate broader geographic coverage.
Geolocation, the task of determining an image’s geographical location, is crucial for applications like crime tracking, navigation, fact-checking, and cultural exploration (Cheng et al., 2022; Chalvatzaras et al., 2022). It involves interpreting contextual clues within an image, such as architectural styles, road signs, natural landscapes, and cultural markers. Inferring location from such diverse indicators demands advanced reasoning, making geolocation a challenging task for both artificial models and human experts (Khan et al., 2024).
Significant effort has been devoted to solving the geolocation task, but often at a coarse level of granularity. For example, methods like Im2GPS3K (Vo et al., 2017a) and PlaNet (Weyand et al., 2016a) frame the task by dividing the globe into grid cells and training deep neural networks to predict the correct cell for a given image. Subsequent studies improve precision by retrieving the most visually similar image from a dataset and using its coordinates as the predicted location (Zhu et al., 2022; Müller-Budack et al., 2018). The reason for this coarse granularity in many approaches is potentially due to the lack of high-quality datasets. For example, Im2GPS3K contains up to 35% non-localizable images (Astruc et al., 2024), while the YFC100M dataset includes irrelevant data such as indoor photos and food images, which provide little to no locational information (Theiner et al., 2022a). Additionally, many datasets are limited in size, with Georeasoner (Li et al., 2024a) featuring only 3K images, thereby restricting the robustness and generalizability of geolocation models. A comparison of these datasets is shown in Table 1.
To address the above obstacles, in this work, we leverage the contributions of hundreds of thousands of geolocation game enthusiasts who provide real user prediction annotations while playing the game. Specifically, we launched a free, public-benefit-oriented online geoguessing platform in June 2022, as shown in Figure 1(a). A screenshot of the platform’s GUI is provided in Appendix A. In each game, two players independently guess the location based on the same image and their own hints, with scores determined by the distance between their predictions and the ground-truth location. The images are sourced from Google Maps, Baidu Maps, Tencent Maps, and Gaode Maps. The platform offers multiple game modes, allowing users to either choose opponents or join random matchups. As of December 17, 2024, this platform has 740,468 users, 3,954,397 locations as unique geolocation tasks, and 25,355,174 human response records. We name the collected dataset GeoComp. This rich and valuable dataset of real human responses enables us to evaluate task difficulty and filter out unreasonable cases. For instance, some tasks are too easy, such as when the name of a shopping mall in a city is clearly visible in the image, enabling most users to answer correctly. On the other hand, some tasks are highly challenging, where only a few users spend considerable time before providing accurate answers. Additionally, there are unreasonable tasks that contain no identifiable hints, making them unsolvable for all users despite significant effort.
Unlike previous approaches that address this task with a coarse level of granularity, we conduct a comprehensive evaluation of recent advanced LVMs on GeoComp, where the models are required to reason and predict the exact city of a given location. Our findings reveal that this task poses a significant challenge for existing LVM models. To address this, we introduce a Geographical Chain of Thought (GeoCoT) approach, which automatically guides the reasoning process through multi-step analysis of geographical cues, such as landmarks, environmental features, and spatial relationships. For the evaluation of the reasoning process, we propose a set of articulated evaluation metrics, named as GeoEval including comparison with ground truth reasoning data and intrinsic evaluation. The results demonstrate that our GeoCoT paradigm significantly improves geolocation accuracy. It not only helps break down complex tasks into manageable reasoning steps but also enhances the interpretability of the inference process.
Our work makes key contributions to geolocation. First, we present GeoComp, a large-scale, human-annotated geolocation dataset with over 3.9 million location images with corresponding location labels, and 25 million human player annotations, featuring diverse geographic regions, languages, and environmental contexts. These annotations identify high-difficulty geolocation cases and establish benchmarks to guide future advancements. Second, we introduce the Geographical Chain of Thought (GeoCoT) framework, a multi-step reasoning approach that improves geolocation accuracy by leveraging geographical cues like landmarks, environmental features, and spatial relationships. Finally, through comprehensive evaluations involving human assessments and LLM inferences, we show that GeoCoT improves predictive performance by up to 25% while enhancing interpretability.
Figure 1. (a) The gaming logic of our platform: Two players independently guess the location based on the same image and their own hints, with scores determined by the distance between their predictions and the ground truth location. (b) The global map shows spatial heterogeneity, with dense clusters in more urbanized regions like Europe and Asia, and sparse coverage in areas like Africa and Siberia. (c) The pie chart highlights the proportional geo-tagged locations distribution, led by North America and Asia. (d) Unlike previous datasets like OSV-5M, where a single country (e.g., America) dominates 25% of the data, our dataset is balanced at country level.
2. Related work
2.1. Image Geolocation Task
Image geolocation refers to determining the corresponding location of a given image, a crucial task in computer vision (Zhu et al., 2023b, a, 2024a), spatial data mining (Zhao et al., 2017, 2016; Zhang et al., 2023d; Han et al., 2023), and GeoAI (Zhao and Tang, 2017; Zhao et al., 2022; Zhang et al., 2023a, e). Previous research in image geolocalization could be primarily classified into two approaches: classification-based methods and retrieval-based methods. (1) Classification-based methods partition most regions of the Earth into multiple grid cells. Models are trained to classify each image into the correct cell (Clark et al., 2023b; Pramanick et al., 2022; Müller-Budack et al., 2018; Seo et al., 2018; Weyand et al., 2016b). The center coordinates of each cell are used as the predicted values. However, due to the limited number of cells, the granularity of the predicted values is coarse, preventing precise predictions. (2) Retrieval-based methods establish a database of geographic images with GPS coordinates. For a given input image, these methods retrieve the most similar image from the dataset and use its coordinates as the predicted location (Zhu et al., 2022; Müller-Budack et al., 2018; Zhang et al., 2023b; Workman et al., 2015; Liu and Li, 2019; Zhou et al., 2024). However, constructing a comprehensive global-level image database is clearly impractical.
2.2. Geolocation Dataset
Existing geolocation datasets primarily originate from web-scraped or street-view images that have not been human-validated, raising concerns about their quality for effectively evaluating geolocation capabilities. For instance, datasets derived from web scraping, such as YFCC100M (Theiner et al., 2022b) and Im2GPS3K (Vo et al., 2017a), include a significant proportion of images depicting food, art, pets, and personal photographs. These types of images are often weakly localizable or entirely non-localizable (Theiner et al., 2022a). Street-view datasets also face limitations, such as restricted geographic coverage (Astruc et al., 2024); for example, (Mirowski et al., 2019)’s work includes data from only three cities in the United States. Furthermore, dataset collection processes often introduce biases. For instance, some commonly used platforms are inaccessible in certain countries, resulting in uneven geographic representation. Additionally, the difficulty of individual geolocation tasks varies widely within these datasets, but this aspect has not been comprehensively evaluated. For example, images taken at prominent landmarks are relatively easy to geolocate, while others offer no clear hints and are highly challenging (Astruc et al., 2024). These limitations undermine the reliability of current geolocation benchmarks.
Figure 2. Performance of game players of different levels in mainstream countries. Experts are defined as the top 15% in performance scores, while beginners are those in the bottom 15%.
2.3. Large Vision Language Models
LLMs have exhibited extraordinary emergent abilities by scaling up data and model sizes, notably including instruction following (Li et al., 2023b; Dai et al., 2024), in-context learning (Brown et al., 2020), and Chain of Thought (CoT) (Kojima et al., 2022). Building on these emergent capabilities, significant research efforts have focused on enhancing cross-modality understanding and reasoning capabilities. Numerous studies have been conducted on various aspects of LVMs, encompassing structural design (Liu et al., 2024a; Cai et al., 2023; Liu et al., 2024c), data construction (LAION, 2023; Zhao et al., 2024), training strategies (Zhao et al., 2023; McKinzie et al., 2024; Lu et al., 2024), evaluations (Bithel and Bedathur, 2023), and the development of lightweight LVMs (Zhu et al., 2024b). Additionally, the robust capabilities of LVMs have been applied to other fields, such as medical image understanding (Zhou and Huang, 2024; Zhang et al., 2023c, 2024) and document parsing (Ye et al., 2023; Liu et al., 2024b). Furthermore, the development of multi-modal agents has advanced real-world applications, including embodied agents (Huang et al., 2023; Peng et al., 2024) and GUI agents (Yang et al., 2023; Li et al., 2024b; Song et al., 2024). However, the reasoning capabilities of LVMs in geolocation tasks remain underexplored. One of the primary reasons for this limitation is the lack of high-reasoning-value geographic data.
3. Data Overview
In this section, we first describe the data collection platform, then present the statistical distributions with visualizations, and finally showcase the performance of human players on the dataset, a unique feature that sets our dataset apart.
3.1. Geolocation Competition
Inspired by geoguessr website, we developed a free geolocation game platform that tracks participants’ competition histories. Unlike most geolocation websites, including Geoguessr, which rely solely on samples from Google Street View, our platform integrates Baidu Maps, Tencent Maps, and Gaode Maps to address coverage gaps in regions like mainland China, ensuring broader global accessibility. Users can choose specific opponents or engage in random matches. Each match consists of multiple questions, and each user is initially assigned a “vitality score.” Users mark their predicted location on a map, and the system evaluates accuracy based on the surface distance between the predicted point and the ground truth. Larger errors result in greater deductions from the user’s vitality score. The user with the higher vitality score at the end of the match is declared the winner. To ensure predictions are human-generated rather than machine-generated, users must register with a phone number, enabling tracking of individual activities. Using this platform, we collected GeoComp, a comprehensive dataset covering 1,000 days of user competition.
3.2. Dataset Statistic
Figure 1(a) presents a global heatmap of geo-tagged locations density, highlighting significant spatial heterogeneity in our dataset. High-density regions are concentrated in urbanized zones such as Europe, North America, and parts of Asia, whereas areas like Africa and Oceania are sparsely represented, often due to underdeveloped infrastructure or low population density. Figure 1(b) provides an overview of the proportional distribution of geo-tagged locations counts across continents, offering a macroscopic view of the dataset’s global spread. In Figure 1(c), we further illustrate the geo-tagged locations distribution by country. Notably, in contrast to datasets like OSV-5M (Astruc et al., 2024), which suffers from severe imbalances—such as the U.S. alone accounting for up to 25% of the total data—our dataset achieves a more balanced global distribution. No single country or continent dominates the dataset, ensuring a more equitable geographic representation and highlighting areas where further data collection efforts may be needed.
3.3. Human Player Performance
Figure 3. Comparison of previous geolocation tasks and our proposed paradigm: while previous works focused on coarse-grained predictions limited by dataset quality, our generation and reasoning-based method enables fine-grained city-level predictions.
Our dataset not only includes image and location information but also rich human player performance data on the task. This label information serves not only as a valuable metric for evaluating the difficulty of different tasks but also as a benchmark for understanding human decision-making in geolocation challenges. In this subsection, we analyze the performance score across players and countries, providing insights into how human players perform on a global scale and how their accuracy varies across different regions and task types. We use GeoGuessr’s scoring formula to evaluate a user’s performance on a single task:
$S=\textstyle\left\lfloor\exp\left(-\frac{d}{s_{d}}\right)\times 5000\right\rfloor.$
Here, $S$ is the user’s score (0 to 5000), $d$ is the geographic distance between the predicted and actual locations (in kilometers), and $s_{d}$ is the maximum distance within the area divided by 10. The score decreases exponentially as $d$ increases. For example, $s_{d}$ is 1805 km globally (based on Earth’s maximum distance of 18,050 km) and 615 km for China, reflecting smaller scales. A perfect prediction ($d=0$) yields $S=5000$. A player’s performance score is defined as the average score across all their tasks. Similarly, a country’s performance score is the average score across all tasks performed within that country.
Player Performance Across Levels. The performance of game players across different levels, as illustrated in Figure 2, highlights significant gaps between beginners and experts in mainstream countries. Expert players, defined as the top 15% of performers, consistently achieve much higher performance metrics compared to beginners, defined as the bottom 15%, with noticeable gaps in countries like Canada, China, and India. For example, in Canada, experts perform nearly 10 times better than beginners, underscoring the steep learning curve involved in mastering the game. This performance gap presents challenges for new players, as it emphasizes the level of skill, strategy, and game knowledge required to compete effectively at higher levels.
Player Performance Across Countries. The player performance across countries, as shown in Figure 2, demonstrates significant variations influenced by three key factors: climate, geographic size, and language. Players tend to perform well in countries such as Germany, France, and Japan. These nations are characterized by unique languages and relatively small geographic sizes. The presence of distinctive languages on urban street signs provides clear linguistic clues, enabling players to quickly identify the country. Additionally, the compactness of these countries allows for more precise guesses, resulting in higher scores. In contrast, despite China’s unique language, its vast size and diverse climates make pinpointing specific locations challenging, leading to lower scores. Similarly, large countries like the USA, China, and Canada face additional challenges due to their shared temperate climates and extensive territories, where players often confuse them due to similar vegetation and climate, reducing accuracy.
Player Performance Across Tasks. From Figure 2, we can also observe significant variations in player performance across different tasks. In certain countries, player performance is relatively low, while in others, it is notably higher. This highlights the diversity in task difficulty within our dataset, offering valuable insights for assessing and categorizing task complexity.
Table 2. Comparison of Precision, Recall and F1 scores in country-level and city-level geolocation. The scores are represented as follows: best, second, and third. Numbers in bold mean that the improvement to the best baseline is statistically significant (a two-tailed paired t-test with p-value <0.01).
Model City Country Continent
[…]
LLaVA-1.6 0.002 0.001 0.002 0.041 0.019 0.028 0.175 0.067 0.056
LLama-3.2-Vision 0.081 0.037 0.030 0.630 0.199 0.217 0.866 0.643 0.639
Qwen-VL 0.016 0.013 0.014 0.069 0.042 0.070 0.130 0.115 0.077
GeoCLIP 0.018 0.007 0.008 0.550 0.197 0.204 0.872 0.746 0.731
GeoReasoner 0.018 0.014 0.012 0.092 0.053 0.085 0.208 0.161 0.144
GPT-4o 0.092 0.048 0.044 0.615 0.188 0.208 0.807 0.468 0.487
GPT-4o(CoT) 0.094 0.052 0.042 0.623 0.194 0.212 0.819 0.430 0.449
GeoCoT 0.118 0.089 0.086 0.640 0.260 0.291 0.862 0.638 0.646
4. Geographic Chain of Thought
In this section, we introduce GeoCoT, a novel chain-of-thought prompting framework for graph-based and geolocation tasks. Unlike standard CoT prompting which performs generic step-by-step reasoning, GeoCoT introduces a domain-specific, hierarchically structured reasoning process that mimics how humans localize geographic information from broad regions to fine-grained details.
4.1. Rethinking Geolocation Task
As discussed in § 2.1, the geolocation task has traditionally relied on classification-based (Clark et al., 2023b; Pramanick et al., 2022; Weyand et al., 2016b) and retrieval-based methods (Zhu et al., 2022; Zhang et al., 2023b), as shown in Figure 3. While these approaches have advanced the field, they face significant limitations in precision and scalability, prompting a rethinking of the task.
Inspired by how humans gradually narrow down locations from broad to fine-grained observations (Luo et al., 2022), we propose a new geolocation paradigm: predicting geographic locations through a step-by-step reasoning process. Unlike traditional approaches limited by grid-based classification and exhaustive databases, our model generates natural language reasoning, guiding it to the final predicted city. To implement this paradigm, we introduce GeoCoT (Geographic Chain-of-Thought), a framework designed for both interpretability and accuracy.
4.2. GeoCoT Deisgn
Our design of GeoCoT is inspired by how humans intuitively approach geolocation—progressing from broad to fine-grained analysis. Rather than relying on generic step-by-step reasoning like standard CoT prompting, GeoCoT mimics the human process: starting with macro-level cues (e.g., climate, terrain), then narrowing down to country, city, and finally micro-level details to guide the model through interpretable geographic reasoning.
Concretely, GeoCoT operates in five sequential stages: 1. Continental or Climate Zone Identification. The process begins with identifying broad regions using natural features like mountains, vegetation, or soil, narrowing the scope to a continent or climate zone. 2. Country-Level Localization. Cultural markers, language on signs, and architectural styles are analyzed to refine predictions to the country level. 3. City-Level Refinement Using Infrastructure. Street elements, such as driving direction, bollards, and license plate colors, are used to locate specific cities or regions. 4. Landmark-Based Verification. Features like fire hydrants, guideposts, and street signs help validate and further refine the predicted location. 5. Fine-Grained Micro-Level Validation. Finally, subtle details such as sidewalk patterns and clothing styles confirm precise localization at a city or neighborhood level. These five reasoning steps are formulated as a single, structured prompt and jointly fed into the LVM, which directly generates the final predicted location. Detailed prompts can be found in Appendix B.
It is important to note that GeoCoT does not require any concrete knowledge about the specific features of locations. Instead, it offers reasoning tutorials designed to help LVMs identify geographic clues by leveraging their existing knowledge.
5. Experiments
In this section, we first introduce our experimental settings, then evaluate GeoCoT in terms of its general geolocalization ability, followed by a detailed evaluation of its reasoning process.
5.1. Setting
We selected 500 geo-tagged locations with high inferential value from the dataset to serve as a test set, using a stratified sampling method across continents to ensure balanced geographic distribution. This number is larger than in previous works (Liu et al., 2023b; Guan et al., 2024), which typically include only a few dozen case studies to examine the reasoning process. We define ”high inferential value” as locations with moderate difficulty—challenging enough to be correctly identified by experienced participants, but not so easy that beginners can do so effortlessly. Specifically, we selected 20 mainstream countries across six continents as representative samples and extracted tasks with an average player score of around 3,000 out of 5,000 for annotation. This test set has been publicly released on GitHub. Our GeoCoT framework is implemented using GPT-4o.
5.2. Baselines
We compare our model, GeoCoT, against several strong baselines representing the latest advancements in geolocation on our dataset. General Open-Source VLMs: LLaVA-1.6 (Liu et al., 2023a) utilizes a fully connected vision-language connector, effectively bridging visual inputs with linguistic features to deliver strong results in geolocation tasks. Llama-3.2-vision (Meta, 2024) demonstrates advanced multi-modal reasoning capabilities, making it a powerful open-source vision-language model. Qwen-VL (Bai et al., 2023), leveraging vast datasets of billions of image-text pairs, achieves robust performance in geolocation through its strong visual and spatial semantic understanding. Baselines Targeting Geolocation Tasks: GeoCLIP (Vivanco Cepeda et al., 2024), inspired by CLIP, aligns images with GPS coordinates using a retrieval-based approach to enhance geolocation. GeoReasoner (Li et al., 2024a) combines geospatial reasoning with visual-language alignment for state-of-the-art geolocation performance. Closed-Source VLMs: GPT-4o (OpenAI, 2024) excels in vision reasoning tasks with its advanced multi-modal capabilities. Furthermore, GPT-4o(CoT), following the setting of cot-zero-shot (Wei et al., 2022), leverages chain-of-thought reasoning to improve performance in complex scenarios. All models are evaluated using the same input format and test set to ensure a fair comparison.
5.3. Overall Performance Evaluation of GeoCoT
In this subsection, we evaluate the city location prediction performance of our model in comparison with the latest LVM models. We evaluate geolocation performance from two aspects: first, location prediction compared with the ground truth at various levels; and second, the direct calculation of the Earth’s surface distance. We present the location prediction performance in Table 2, evaluated across three levels: city, country, and continent. Performance is measured using accuracy, which calculates the proportion of correct predictions out of all predictions; recall, which determines the proportion of true positive predictions out of all actual positive cases; and the F1 score, which balances precision and recall to provide their harmonic mean.
The results reveal several key observations. First, open-source LVMs such as LLaMA-3.2-Vision achieve competitive performance, performing on par with GPT-4o and GPT-4o (CoT), demonstrating their effectiveness in location prediction tasks. Second, performance varies across different levels of granularity. While GPT-4o (CoT) ranks second at the city level, it underperforms at the country level, highlighting the importance of multi-level evaluation to fully assess a model’s geolocation reasoning ability. Finally, our model, GeoCoT, consistently achieves top performance across all nine metrics and three levels, demonstrating its robustness and adaptability in geolocation tasks. Additionally, GeoCLIP surpasses GPT-4o at the continent level, which can be attributed to its pretraining on image-GPS pairs, making it particularly well-suited for coarse-grained geolocation tasks. Coarse-grained continent-level predictions typically require less detailed local knowledge and instead rely on broader geographic cues, such as climate, landscapes, and cultural markers. However, GeoCLIP performs poorly at finer granularities like country and city levels, suggesting that it lacks a strong capability for geographic reasoning beyond direct visual features.
Table 3. Accuracy of different models on geolocation tasks at various scales. Numbers in bold mean that the improvement to the best baseline is statistically significant (a two-tailed paired t-test with p-value <0.01).
Model Street City Country
1km 25km 750km
LLaVA-1.6 0.006 0.020 0.082
Llama-3.2-Vision 0.018 0.104 0.638
Qwen-VL 0.004 0.014 0.090
GeoCLIP 0.035 0.077 0.625
GeoReasoner 0.010 0.020 0.128
GPT-4o 0.045 0.147 0.678
GPT-4o(CoT) 0.047 0.151 0.701
GeoCoT 0.073 0.157 0.711
Next, in Table 3, we present the accuracy of each model by measuring the geographic distance between the predicted city and the ground truth. The metrics represent the proportion of predictions within three distance thresholds: Street (1 km), City (25 km), and Country (750 km). Higher values indicate better performance, with stricter thresholds assessing fine-grained localization and larger thresholds evaluating coarse-level accuracy. The results show that GPT-4o and Llama-3.2-vision outperform the dedicated large-scale model GeoCLIP for geolocation, even under finer-grained evaluation settings. For example, at the street-level threshold, GPT-4o achieves 0.045 compared to GeoCLIP’s 0.035, and at the city-level threshold, GPT-4o scores 0.147, nearly double GeoCLIP’s 0.077. Moreover, our proposed GeoCoT paradigm demonstrates even greater improvements. At the street level, GeoCoT achieves 0.073, significantly outperforming both GeoCLIP (0.035) and GPT-4o (0.045). Similarly, at the city level, GeoCoT achieves 0.157, and at the country level (750 km), it achieves 0.711, the highest among all models. These results highlight GeoCoT’s strong performance and the potential of its reasoning framework for geolocation tasks.
5.4. GeoEval: Reference-Based Evaluation of GeoCoT Reasoning
Beyond evaluating overall task performance, we focus on analyzing the reasoning process of GeoCoT, which emulates a human-like reasoning approach. To establish a reference for this evaluation, three gaming enthusiasts collaboratively constructed reasoning processes for the same 500 cases based on geo-tagged locations. We designated these as the reasoning ground truth (a human-annotated example can be found in Appendix C). These GT annotations serve as a benchmark within our evaluation framework, GeoEval. The evaluation process utilizes (1) GPT-based assessment through GPTScore (Fu et al., 2023) and (2) prompt-based scoring.
Our prompt-based scoring includes four dimensions ranging from 0-5, and the detailed prompts can be found in Github. The first dimension is the completeness of feature extraction (CE), which evaluates whether all key clues provided in the GT are comprehensively covered and accurately described in the reasoning process. Comprehensive feature extraction ensures that reasoning outcomes are based on sufficient factual evidence, thereby enhancing their reliability and accuracy. The second dimension is the accuracy of feature extraction (AE), which measures whether the identified and described attributes or characteristics of the key information in the GT are correct. Misidentified features can lead to reasoning outcomes that deviate from the facts, reducing the credibility of the results. The third dimension is the accuracy of reasoning and cue correspondence (AC), which assesses whether the reasoning process derives reasonable conclusions based on the extracted cues and maintains consistency with the reasoning logic presented in the GT. Incorrect correspondence between cues and conclusions can result in outcomes that deviate from reality. The final dimension is the logical coherence of reasoning (LC), which evaluates the consistency, logical flow, and adherence to common sense within the reasoning chain. Logical errors compromise the reliability of the reasoning process and hinder the model’s ability to arrive at accurate conclusions.
Table 4. Evaluation of GeoCoT’s reasoning process using ground truth-based metrics within the GeoEval framework. Numbers in bold mean that the improvement to the best baseline is statistically significant (a two-tailed paired t-test with p-value <0.01).
Model Similarity GeoEval
GPTScore CE AE AC LC
LLaVA-1.6 0.478 1.262 1.271 1.446 1.490
Llama-3.2-Vision 0.566 2.203 2.386 2.558 2.721
Qwen-VL 0.371 1.231 1.255 1.453 1.484
GeoReasoner 0.424 1.421 1.533 1.719 2.038
GPT-4o 0.613 2.320 2.891 2.809 3.143
GPT-4o(CoT) 0.663 2.462 3.136 3.156 3.540
GeoCoT 0.728 2.690 3.538 3.696 3.945
The experimental results in Table 4 highlight the significant advantages of GeoCoT compared to baseline models across all evaluation metrics. GeoCoT achieves the highest GPTScore of 0.728, outperforming GPT-4o (CoT) (0.663) and -1.6 (0.478), demonstrating its superior alignment with human-constructed reasoning processes. In terms of feature extraction, GeoCoT achieves a CE score of 2.690 and an AE score of 3.538, significantly surpassing GPT-4o (CoT) and the dedicated GeoReasoner model. Furthermore, GeoCoT’s performance in reasoning accuracy and logical coherence is unmatched, with AC and LC scores of 3.696 and 3.945, compared to GPT-4o (CoT), which scores 3.156 and 3.540, and GeoReasoner, which lags behind at 1.719 and 2.038. These results clearly demonstrate that GeoCoT not only captures key information more comprehensively but also maintains a more accurate and logically coherent reasoning process compared to both reasoning-based models and traditional baselines like GeoReasoner.
5.5. Intrinsic Evaluation of GeoCoT Reasoning
Figure 4. Qualitative comparison of LLaVA, GPT4o, and GeoReasoner. Clues are shown in blue, correct predictions in green, incorrect in red, and vague/uncertain guesses in orange.
We begin with a ground truth-based evaluation, comparing GeoCoT’s reasoning to human-authored processes to assess its alignment with established reasoning patterns. To complement this, we conduct an intrinsic evaluation focused on hallucination errors—assessing logical consistency, coherence, and robustness without relying on external references. Given the need for multimodal judgment, this evaluation is performed by human annotators.
Following previous work on assessing hallucinations in terms of objects, attributes, and relationships (Li et al., 2023a; Sun et al., 2023), we evaluate the quality of synthetic data across three key dimensions: (1) Object Hallucination (OH): assesses whether the synthetic data includes objects that do not exist in the image. Object Hallucination evaluates the extent to which synthetic data introduces fictional elements. (2) Fact Hallucination (FH) measures the accuracy of factual information within the synthetic data. Fact Hallucination occurs when the synthetic data contains facts, figures, or other information that is incorrect or not supported by the original data. (3) Attribution Hallucination (AH) evaluates whether the synthetic data incorrectly attributes properties, characteristics, or relations to entities or objects. To quantify hallucinations, each detected error is counted as one instance in the corresponding dimension. To evaluate these dimensions, we invited 2 human annotators with professional backgrounds in geographic reasoning and data validation to assess GPT-4o, GeoReasoner, and our proposed GeoCoT model. These three baselines provide textual reasoning processes across 1,500 evaluated cases. The results, shown in Table 5, indicate the number of errors in each dimension, demonstrating that GeoCoT significantly reduces hallucination errors compared to the other models. The inter-annotator agreement, measured by Cohen’s Kappa, is 0.82 for OH, 0.79 for FH, and 0.85 for AH, indicating substantial agreement across all dimensions. The correlation between the human evaluation scores and the automatic evaluation metrics in Table 2 is –0.99 (p ¡ 0.01), demonstrating a strong inverse relationship: as hallucination errors decrease, overall geolocation performance improves significantly.
Table 5. Hallucination Evaluation on Reasoning Data.
Model OH FH AH
[…]
GeoReasoner 237 151 203
GPT-4o 43 4 35
GeoCoT 5 1 18
Table 6. Performance comparison of GeoCoT and state-of-the-art geolocation models on traditional benchmarks. Numbers in bold mean that the improvement to the best baseline is statistically significant (a two-tailed paired t-test with p-value <0.01).
Model Im2GPS Im2GPS3K
Street City Country Street City Country
1km 25km 750km 1km 25km 750km
LLaVA-1.6 0.04 0.18 0.39 0.03 0.14 0.32
Llama-3.2-Vision 0.09 0.37 0.65 0.07 0.27 0.52
Qwen-VL 0.04 0.21 0.37 0.04 0.15 0.26
GeoCLIP 0.17 0.41 0.77 0.13 0.32 0.67
GeoReasoner 0.05 0.19 0.33 0.04 0.15 0.26
GPT-4o 0.13 0.47 0.74 0.14 0.40 0.66
GPT-4o(CoT) 0.16 0.49 0.77 0.14 0.45 0.69
GeoCoT 0.22 0.55 0.83 0.15 0.46 0.74
5.6. Case Study
We present two examples in Figure 4 to analyze the performance of LLaVA, GPT4o, and GeoReasoner, highlighting the effectiveness of our GeoCoT approach. In the first example, struggles to provide a specific prediction, reflecting its reliance on general architectural cues and its tendency to consider broad regions such as the United Kingdom or France. GPT4o, despite identifying key features of the European landscape, incorrectly associates them with Germany, indicating limitations in handling specific regional markers. In contrast, GeoCoT accurately pinpoints the location in France by effectively integrating textual clues, architectural elements, and environmental context.
In the second example, GeoCoT correctly identifies the location as San Francisco, USA, by analyzing U.S. traffic standards, license plates, and local signage, demonstrating strong contextual reasoning. LLaVA-1.6 makes a broad prediction, covering the U.S., Australia, and the U.K., showing uncertainty from general cues. GPT-4o misidentifies the scene as Seattle, relying on architectural similarities but missing key details.
5.7. Generalizability Evaluation
Even though our dataset is more comprehensive and human-annotated, we are also interested in evaluating how our model performs on traditional geolocation datasets to provide a more thorough comparison. Hence, we select two existing benchmark datasets, Im2GPS (Hays and Efros, 2008) and Im2GPS3K (Vo et al., 2017b), due to their popularity and widespread use in geolocation tasks as standard benchmarks for evaluating model performance. Similarly, we use the center point coordinates of the city text address in GeoCoT’s output and measure the distance between the output and the ground truth locations.
We present the performance results in Table 6. We observe that state-of-the-art geolocation models, such as GeoCLIP, perform well on traditional geolocation tasks, surpassing GPT-4o and coming close to our model, GeoCoT. However, this is in contrast to the results shown in Table 2, where GeoCLIP significantly underperforms GPT-4o on fine-grained city- and country-level geolocation tasks. This discrepancy suggests that these baseline models may be overfitting to the specific datasets they were trained on, lacking the generalization ability required for more diverse or fine-grained geolocation challenges. In contrast, our model consistently outperforms traditional methods across different granularity levels and datasets without any training, and thus does not suffer from overfitting.
6. Conclusion
In this work, we present the largest geolocation dataset to date, collected from a geolocation game platform with 740K users over two years. The dataset comprises 25M entries of metadata, including 3M geo-tagged locations spanning most of the globe, each annotated thousands to tens of thousands of times by human users. This dataset enables diverse difficulty-level analysis and highlights the limitations of current LVMs. We also introduce a generation-based reasoning solution for the geolocation task, where the LVM generates reasoning chains by leveraging clues from images and produces the final predicted location. Using our GeoEval set of metrics, we demonstrate that our GeoCoT framework significantly outperforms state-of-the-art general and task-specific baselines on this dataset.
In future work, we plan to enhance model interpretability and robustness, explore multi-modal integration of text and visuals, and expand the dataset to better cover underrepresented regions for improved global coverage and fairness.
Data Ethics
The creation and release of our dataset adhere to stringent ethical standards to ensure the privacy and well-being of all contributors. We have conducted rigorous anonymization of the dataset to protect user privacy. All personally identifiable information, such as usernames, email addresses, and IP addresses, has been permanently removed. Only non-identifiable behavioral data, such as prediction outcomes and timestamps, are retained. The dataset originates from user participation on our open-source geolocation game platform. Users were informed during the registration process that their activity data might be used for research purposes. This ensures transparency in data collection and maintains user trust. We have explicitly designed the dataset for research purposes, with the sole intention of advancing geolocation and related artificial intelligence technologies. Importantly, our dataset does not include the images directly but instead provides links to images hosted on platforms such as Google Maps or Baidu Maps, which can be accessed through their official APIs.
We are committed to ensuring the responsible use of this dataset. Researchers accessing the data must agree to a data usage agreement that prohibits unethical or illegal use.